An n year deferred perpetuity due has the following payment pattern.

300; 100; 400; 300; 100; 400; 300; 100; 400; 300; 100; 400; ::::

The perpetuity has a price of 694.5262387, and this price is based on an annual effective interest rate of 10%. The payments occur on an annual basis. Find n
Answer #1

694.5262387=(300/1.1^(n)+100/1.1^(n+1)+400/1.1^(n+2))*1/(1-1/1.1^3)
=>694.5262387*(1-1/1.1^3)=1/1.1^(n-1)*(300/1.1+100/1.1^2+400/1.1^3)
=>1/1.1^(n-1)=694.5262387*(1-1/1.1^3)/(300/1.1+100/1.1^2+400/1.1^3)
=>1/1.1^(n-1)=0.26333

=>1.1^(n-1)=1/0.26333
=>(n-1)*log(1.1)=log(1/0.26333)
=>n=1+log(1/0.26333)/log(1.1)
=>n=15.00005